This is a special order problem that also requires that you use the high low method to estimate some cost function parameters, so you may want to review the high-low method lectures in Module 1. As with almost all of the analyses that we have done, determining variable and fixed costs, and knowing what to do with them, is critical.
______________________________________________________


Huang Automotive is presently operating at 75% of capacity. The company recently received an offer from a Korean truck manufacturer to purchase 28,500 units of a power steering system component for $195 per unit.  Peter Wu, vice-president of sales, notes that although there will be an additional $2.00 shipping cost for each component, he thinks that accepting the order will get the company's "foot in the door" of an expanding international market.

To determine variable and fixed costs, Huang's accountant used the high-low method with the following production and cost information for the last two years:

196,000 units     233,000 units    
Direct material costs $17,052,000      $20,271,000     
Direct labor costs 4,900,000      5,825,000     
Overhead costs 22,172,000      24,281,000     
Selling and administrative costs 12,054,000      12,479,500     
Total costs $56,178,000      $62,856,500     
Total costs per unit $286.62      $269.77     



T.J. Chan, vice-president of engineering, feels that any new market should first show its profitability and that the $195 per unit offer is not only below the regular $260 selling price, but it's below the unit cost of the component. She also points out that there will be additional setup costs of $225,000 and that Huang will have to lease some special equipment for $270,000.

REQUIRED [6 tries]
1. Using the high-low method to determine cost behavior, what would the expected profit be on the special order (use a negative sign for a loss)?
Answer #1

FIXED COSTS are constant and they will not change with acceptance of order

so they are not relevant for decision making

we will find out fixed and variable cost.

activity LOW a activity HIGH b change a-b variable cost [change in cost/change in activity]
activity 196000 233000 37000
Direct material costs $17,052,000      $20,271,000      3219000 87$[3219000/37000]
Direct labor costs 4,900,000      5,825,000      925000 25$[925000/37000]
Overhead costs 22,172,000      24,281,000      2109000 57
Selling and administrative costs 12,054,000      12,479,500      425500 11.5

material cost per unit =$87*196000units = 17052000$ WHICH IS EQUAL TO TOTAL COST

so whole cost is variable 87$

labor cost = 25*196000=4900000

so whole cost is variable $25

overhead cost = 196000*57

=11172000

total cost = fixed cost+variable cost

22172000= fixed + 11172000

fixed cost =22172000-11172000

=11000000

selling and administration = 11.5*196000=2254000

12054000 =fixed + 2254000

fixed = 9800000$

cost for 28500unitS

SPECIAL ORDER PROFIT

Revenue 5557500 [195*28500]
direct material (2479500) [87*28500]
labor (712500) [25*28500]
overhead [variable part only] (1624500) [57*28500]
selling and administrative [variable part only] (327750) [11.5*28500]
Additional shipping cost (57000) [2*28500]
set up cost (225000)
special equipment (270000)
net income(Loss) -138750

Loss = 138750$